8.7 A business places substantial emphasis on customer satisfaction and, to this end, delivers its
product in special protective containers. These containers have been made in a department
within the business. Management has recently become concerned that this internal supply of
containers is very expensive. As a result, outside suppliers have been invited to submit tenders
for the provision of these containers. A quote of £250,000 a year has been received for a volume
that compares with current internal supply.
An investigation into the internal costs of container manufacture has been undertaken and
the following emerges:
(a) The annual cost of material is £120,000, according to the stores’ records, maintained at
actual historic cost. Three-quarters (by cost) of this represents material that is regularly
stocked and replenished. The remaining 25 per cent of the material cost is a special foaming
chemical. This chemical is not used, by the business, for any other purpose than making
the containers. There are 40 tonnes of this chemical currently held. It was bought in bulk for
£750 a tonne. Today’s replacement price for this material is £1,050 a tonne but it is unlikely
that the business could realise more than £600 a tonne if it had to be disposed of owing to
the high handling costs and special transport facilities required.
(b) The annual labour cost is £80,000 for this department. Most of this cost, however, relates
to casual employees or recent starters. If an outside quote were accepted, therefore, little
redundancy would be payable. There are, however, two long-serving employees who would
each accept as a salary £15,000 a year until they reached retirement age in two years’ time.
(c) The department manager has a salary of £30,000 a year. The closure of this department
would release him to take over another department for which a vacancy is about to be
advertised. The salary, status and prospects are similar.
(d) A rental charge of £9,750 a year, based on floor area, is allocated to the containers department.
If the department were closed, the floor space released would be used for warehousing
and, as a result, the business would give up the tenancy of an existing warehouse for
which it is paying £15,750 a year.
(e) The plant cost £162,000 when it was bought five years ago. Its market value now is £28,000
and it could continue for another two years, at which time its market value would have fallen
to zero. (The plant depreciates evenly over time.)
(f) Annual plant maintenance costs are £9,900 and allocated general administrative costs are
£33,750 for the coming year.
Required:
Calculate the annual cost of manufacturing containers for comparison with the quote using
relevant figures for establishing the cost or benefit of accepting the quote. Indicate any assumptions
or qualifications you wish to make.
Answer #1

Answer:    (Figures in Pound)

We need to do 2 things here:   

   (i) Deriving current annual cost of the containers

   (ii) Deriving the net cost of the Quote, if the company stops manufacturing containers.

(i) Annual Cost of the Containers:

Raw Material + Labour (Variable Cost only)+ Rent + Plant Maintenance Cost

Hence,

=120000+50000+9750+9900 = 189650

(ii) Net Cost of the Quote:

Quote Price- Replacement Price of the Raw Material - Rent that will be earned - Current Market Value of Plant

= 250000-600*40-15750-28000 = 182250

Net Benefit on Outsoucing - 189650-182250 = 7400

Note:

(i) Allocated administrative overheads are not considered because they are fixed expenses for the company.

(ii) It is assumed that on closure Plant will be sold at its current market value i.e. 28000

(iii) In labour department both the labourers are going to remain in the company for next 2 years, hence fixed expense currently from the view point of company.

(iv) Here manager's salary is fixed beacuse he will be transferred in different department.

(v) Variable Costs are considered in decision making as fixed costs are irrelevent frrom decision point of view.